The perimeter of a rectangle varies jointly with its length and width. If the perimeter of a rectangle is 24 inches with a length of 5 and a width of 7, what is a rectangle’s width with a perimeter of 32 inches and a length of 5?
Answer by addingup(3677)

11 is the width:
Perimeter= 2w+2l, let me turn the problem around:
2w+2l= Perimeter. We know the length is 5 and the perimeter 32:
2w+2(5)= 32
2w+10= 32 subtract 10 on both sides:
2w= 22 divide both sides by 2
w= 11